A quality control inspector has drawn a sample of 1919 light bulbs from a recent production lot. If the number of defective bulbs is 11 or more, the lot fails inspection. Suppose 10%10% of the bulbs in the lot are defective. What is the probability that the lot will fail inspection? Round your answer to four decimal places.
Answer #1

X ~ Bin ( n , p)

Where n = 19 , p = 0.10

We have to calculate P(X >= 11) = ?

P(X >= 11) = 1 - P(X <= 10)

= 1 - BINOM.DIST ( number_s , trials, probability_s , cumulative)

= 1 - BINOM.DIST ( 10 , 19 , 0.10 , TRUE)

= 1 - 1

= 0.0000